Suppose your firm is considering investing in a project with the cash flows shown below, that the required rate of return on projects of this risk class is 7 percent, and that the maximum allowable payback and discounted payback statistics for the project are 2.5 and 3.5 years, respectively.

Time:   0   1   2   3   4   5   6
Cash flow:   –$4,800   $1,140   $2,340   $1,540   $1,540   $1,340   $1,140

Use the payback decision rule to evaluate this project.
Answer #1
Payback is the time within which cost of project is recovered back.
Calculation of Payback Period:
Year Cash Flow Cumulative Cash Flow
0 $       -4,800 $   -4,800
1             1,140       -3,660
2             2,340       -1,320
3             1,540            220
4             1,540         1,760
5             1,340         3,100
6             1,140         4,240
Payback period = 2+(1320/1540)
= 2.9 Years
Calculation of discounted payback period:
Year Cash Flow Discount factor Discounted Cash Flow Cumulative Discounted Cash Flow
a b c=1.07^-a d=b*c e
0 $       -4,800      1.0000        -4,800       -4,800
1             1,140      0.9346          1,065       -3,735
2             2,340      0.8734          2,044       -1,691
3             1,540      0.8163          1,257           -434
4             1,540      0.7629          1,175            741
5             1,340      0.7130              955         1,697
6             1,140      0.6663              760         2,456
Discounted Payback = 3+(434/1175)
=               3.4 Years
On the basis of payback decision, project is not accepted as payback period of 2.9 years is beyond the deadline of 2.5 years.
On the basis of discounted payback decision, project is accepted as payback period of 3.4 years is within the deadline of 3.5 years.